questions
(1) If the probability of an event NOT happening is 0.06 what is the probability of it happening?
(2)Evaluate 10C3?
(3)if a teacher grades distribution is over 3 years what are empirical probability for some to get A,C,D when A=39, C=258, and D=124?
(4)If there are 984 flower which 632 are white and 352 are purple what is the probability that you will pick a white flower?

Answer by Boreal(15235)

The probabilities have to add to 1
If one is 0.06, the other has to be (1-0.06)=0.94
10C3=
10!/7!3!
10*9*8*7!/7!*3*2*1
The 7! cancel
10*9*8/3*2*1
720/6=120
=======
Three are 421 outcomes
A=39/421 =0.0926
C=258/421=0.6128
D=124/421=0.2945
That adds up to 0.9999.
======
632/(632+352)
632/984
0.6423